How do I find the standard form of the hyperbola from the equation x^2+6x-y+7=0?
Answer by lwsshak3(11628)

How do I find the standard form of the hyperbola from the equation
x^2+6x-y+7=0?
complete the square
x^2+6x-y+7=0
(x^2+6x+9)-y+7-9=0
(x+3)^2-y-2=0
y=(x+3)^2-2
This is an equation of a parabola with vertex at (-3,-2) of the standard form: y=(x-h)^2+k, (h,k)= (x,y) coordinates of the vertex.
Standard form of equation for a hyperbola: (x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of center. Given equation is a parabola, not a hyperbola.